# arXiv:2511.11225v2[cs.IT] 27 Feb 2026

### v2

@@ IV-B Beampattern Analysis
We now focus on the beampattern, which is defined as the electric field strength received by a $y$-polarized receiver in any given far-field direction. For any direction vector $\bm{\kappa}=\kappa_{0}\mathbf{k}(\theta,\phi)$ and any given beamformer $w(\mathbf{s})$, the beampattern is given by
$P(\bm{\kappa})=\left|\int_{\mathcal{S}}\beta(\bm{\kappa})w(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}\right|=\left|\beta(\bm{\kappa})\widetilde{W}(\bm{\kappa})\right|.$ (71)
Here, $\beta(\bm{\kappa})$ is defined in (25) and $\widetilde{W}(\bm{\kappa})$ is the windowed 2D Fourier transform of $w(\mathbf{s})$, defined as
$\widetilde{W}(\bm{\kappa})=\mathcal{F}_{\mathcal{S}}\{w\}(\bm{\kappa})=\int_{\mathcal{S}}w(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}.$ (72)
@@ IV-B1 Uncoupled CAPA
For uncoupled CAPAs, the optimal beamformer is the matched filtering solution; $i.e.,$ $w_{\mathrm{opt}}(\mathbf{s})=\gamma h^{*}(\mathbf{s})$, where $\gamma$ is a scaling factor to satisfy the uncoupled power constraint discussed in Remark 2. Following (III-B2), $\widetilde{W}(\bm{\kappa})$ for uncoupled CAPAs can be calculated as
$\widetilde{W}_{\text{uncoupled}}(\bm{\kappa})$ $=\gamma\int_{\mathcal{S}}h^{*}(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}$
$\hskip-28.45274pt=\gamma\beta^{*}(\bm{\kappa}_{\mathrm{r}})L_{x}L_{y}\mathrm{sinc}\left(\frac{\Delta\bar{\kappa}_{x}L_{x}}{2}\right)\mathrm{sinc}\left(\frac{\Delta\bar{\kappa}_{y}L_{y}}{2}\right)$
$\hskip-28.45274pt=\gamma H(\bm{\kappa}),$ (73)
where $\Delta\bar{\kappa}{x}$ and $\Delta\bar{\kappa}_{y}$ are the $x$ and $y$ components of $(\bm{\kappa}-\bm{\kappa}_{\mathrm{r}})$, and $H(\bm{\kappa})$ is essentially the windowed 2D Fourier transform of $h^{*}(\mathbf{s})$. Consequently, the uncoupled beampattern is
$P_{\text{uncoupled}}(\bm{\kappa})=\big|\gamma\beta(\bm{\kappa})H(\bm{\kappa})\big|.$ (74)
@@ IV-B2 Coupled CAPA
Based on Theorem 1, the optimal beamformer for a coupled CAPA is proportional to $v(\mathbf{s})$; $i.e.,$ $w_{\mathrm{opt}}(\mathbf{s})=\widetilde{\gamma}v(\mathbf{s})$, where $\widetilde{\gamma}$ is a scaling factor to fulfill the coupled power constraint. The integral equation (28) is a spatial convolution over the finite aperture $\mathcal{S}$. By assuming the aperture is large, the convolution theorem can be applied to approximate this relationship in the wavenumber domain as
$C(\bm{\kappa})\cdot\mathcal{F}_{\mathcal{S}}(v)(\bm{\kappa})\approx H(\bm{\kappa}),$ (75)
where $C(\bm{\kappa})=Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})$ denotes the Fourier transform of the overall coupling kernel $c(\mathbf{s})$. It is important to note that this approximation is accurate only when the transmit aperture is sufficiently large. For practical finite apertures, this model ignores significant windowing effects. Based on this large-aperture approximation, $\widetilde{W}(\bm{\kappa})$ for coupled CAPAs is approximated by
$\widetilde{W}_{\text{coupled}}(\bm{\kappa})=\widetilde{\gamma}\mathcal{F}_{\mathcal{S}}(v)(\bm{\kappa})\approx\frac{\widetilde{\gamma}H(\bm{\kappa})}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})}.$ (76)
@@
$P_{\text{coupled}}(\bm{\kappa})\approx\left|\frac{\widetilde{\gamma}\beta(\bm{\kappa})H(\bm{\kappa})}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})}\right|.$ (77)
@@ IV-B3 Impact of Mutual Coupling
$S(\bm{\kappa})=\frac{P_{\text{coupled}}(\bm{\kappa})}{P_{\text{uncoupled}}(\bm{\kappa})}\approx\frac{1}{\varrho\left(Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})\right)},$ (78)
@@ V Extension to Spatially Discrete Arrays
An SPDA model can be obtained from the proposed model of CAPAs. We model the SPDA by partitioning the continuous surface $\mathcal{S}$ into $N$ uniform, non-overlapping antenna elements. Let $\mathcal{S}_{n}\subset\mathcal{S}$ denote the surface of the $n$-th antenna, centered at position $\mathbf{p}_{n}$. We assume that these element surfaces are disjoint; $i.e.,$ $\mathcal{S}_{n}\cap\mathcal{S}_{m}=\emptyset$ for $n\neq m$, and each has an identical area $|\mathcal{S}_{n}|=A_{\mathrm{d}}$. Furthermore, we first define a reference element subsurface $\mathcal{S}_{\mathrm{d}}$ centered at the origin as
@@
where $L_{\mathrm{d},x}L_{\mathrm{d},y}=A_{d}$. The surface of the $n$-th antenna is then a simple translation of this reference; $i.e.,$ $\mathcal{S}_{n}=\mathcal{S}_{\mathrm{d}}+\mathbf{p}_{n}$.
@@
where $v_{n}\in\mathbb{C}$ is the discrete beamforming coefficient for the $n$-th antenna, $a_{\mathrm{t}}(\mathbf{s}-\mathbf{p}_{n})$ is the element’s current profile, shifted to its center $\mathbf{p}_{n}$ and $\Pi_{n}(\mathbf{s})$ is a rectangular window function that confines the current to the $n$-th element’s surface, defined as $\Pi_{n}(\mathbf{s})=1$ for $\mathbf{s}\in\mathcal{S}_{n}$ and $\Pi_{n}(\mathbf{s})=0$ otherwise.
@@ V-A Mutual Coupling
where $\Delta\mathbf{p}_{nm}=\mathbf{p}_{n}-\mathbf{p}_{m}$ denotes the spacing between the $n$-th and the $m$-th antenna elements. Note that the choice of current profile $a_{\mathrm{t}}(\mathbf{s})$ significantly affects the mutual coupling, and it is possible to design $a_{\mathrm{t}}(\mathbf{s})$ carefully to minimize the mutual coupling between different elements. Moreover, the model in (83) is consistent with the mutual-coupling impedance matrices commonly used in circuit-based analyses of SPDAs, e.g., [10547020, Eq. (5)].
@@
Based on the analysis in Proposition 1, a half-wavelength spacing cannot null the mutual coupling for polarized antennas; $i.e.,$ $c_{\mathrm{rad}}(\Delta\mathbf{p}_{nm})\neq 0$ for $\Delta\mathbf{p}_{nm}=\pi/2$. This is in contrast with the case of ideal isotropic antennas, which are widely assumed in the literature.
@@
Figure 3: Comparison of the true radiation mutual coupling kernel (15) with the proposed approximation (35) for various Gauss-Legendre quadrature orders $M$.
@@ V-B Beamforming
where $\bm{h}=[h_{1},\dots,h_{N}]^{H}$ is the channel vector for the SPDA. Consequently, the beamforming optimization problem for the SPDA can be formulated as
@@ VI Numerical Results
This section presents numerical examples to validate the proposed mutual coupling kernel approximation and the corresponding beamforming designs. Unless specified otherwise, all simulations share the following setup. Specifically, the signal frequency is $f=2.4$ GHz. The transmit surface is a square with dimensions $L_{x}=L_{y}=0.5$ m. The receiver is placed in the far-field at a distance of $R_{0}=50$ m, which satisfies the far-field condition $R_{0}\geq 2D^{2}/\lambda=26$ m, with $D=\sqrt{L_{x}^{2}+L_{y}^{2}}$ denoting the aperture of the transmit surface. The transmit surface is modeled as copper, with conductivity $\sigma_{s}=5.8\times 10^{7}$ S/m and permeability $\mu_{s}=4\pi\times 10^{-7}$ H/m. The Gauss-Legendre order is set to $M=20$. For SPDAs, we assume a simple uniform current profile $a_{\mathrm{t}}(\mathbf{s})=1/\sqrt{A_{\mathrm{d}}}$, an element aperture $0.1\lambda\times 0.1\lambda$ [9906802, wong2014design], and an antenna spacing of $\lambda/2$.
Figure 4: Convergence of the optimal array gain with respect to the Gauss-Legendre quadrature orders $M$.
Figure 5: Convergence of the conjugate gradient.
VI-A Convergence of Proposed Methods
We first evaluate the convergence performance of the proposed kernel approximation (KA) and conjugate gradient (CG) methods. First, the accuracy of both methods is highly dependent on the order of the Gauss-Legendre quadrature. In particular, the kernel approximation method exploits the Gauss-Legendre quadrature to approximate the radiation mutual coupling kernel, c.f. (35). Fig. 3 illustrates the true kernel and the approximated kernel under different Gauss-Legendre quadrature orders $M$ at 7.8 GHz. It can be observed that as $M$ increases, the approximated kernel gradually converges to the true kernel, validating the effectiveness of the proposed method. Furthermore, Fig. 5 demonstrates the convergence of the array gain achieved by both methods with respect to $M$. Both methods eventually converge to the same value under different system parameters, which cross-validates the correctness of the results. We note that the CG method converges more smoothly. This may be because KA method exploits the model approximation; $i.e.,$ approximating the coupling kernel itself, whereas the CG method applies numerical approximation; $i.e.,$ approximating the final numerical implementation. Finally, Fig. 5 shows the convergence of the CG method as iterations proceed. As the frequency increases, the CG method requires more iterations to converge. In contrast, the KA method is non-iterative, thus exhibiting significantly lower computational complexity. Given this advantage, the KA method is used to generate the numerical results in the following sections.
@@ VI-B Array Gain
We now evaluate the array gains under different setups, using both coupled and uncoupled SPADs as benchmarks. The uncoupled benchmark is obtained by modifying the matrix $\mathbf{\Psi}$ defined in (V-A), keeping only its diagonal elements and setting all off-diagonal elements to zero.
Fig. 7 depicts the optimal array gain versus the antenna spacing of SPDAs, where smaller spacing implies more antennas are deployed within a fixed aperture. As expected, when mutual coupling is considered, the SPDA array gain gradually converges to the CAPA performance as antenna spacing shrinks. This matches the physical expectation that CAPA is the ultimate limit of the SPDA. Nevertheless, the uncoupled model fails to capture the correct behavior. It underestimates performance at large spacing and significantly overestimates it at small spacing. More partially, the array gain for the uncoupled SPDAs tends to increase unboundedly as spacing reduces, which is a violation of physics. Fig. 7 further shows the array gain versus the aperture size $|\mathcal{S}|$. For coupled CAPAs, the array gain grows linearly as the aperture size increases. However, the SPDA array gain exhibits a stepwise increase, since a larger aperture does not always accommodate more antenna elements.
@@ VII Conclusions
This paper has primarily focused on unpolarized and single-user CAPA systems. Several promising future research directions can be identified as follows:

### v3

@@ IV-B Beampattern Analysis
We now focus on the beampattern, which is defined as the electric field strength received by a $y$-polarized receiver in any given far-field direction. For any wavevector $\bm{\kappa}=\kappa_{0}\mathbf{d}(\theta,\phi)$ and any given beamformer $w(\mathbf{s})$, the beampattern is given by
$B(\bm{\kappa})=\left|\int_{\mathcal{S}}\hat{\beta}(\bm{\kappa})w(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}\right|=\left|\hat{\beta}(\bm{\kappa})\hat{W}(\bm{\kappa})\right|.$ (73)
Here, $\hat{\beta}(\bm{\kappa})=\left(1-\sin^{2}\theta\sin^{2}\phi\right)$ the polarization projection factor as in (25) and $\hat{W}(\bm{\kappa})$ is the windowed 2D Fourier transform of $w(\mathbf{s})$, defined as
$\hat{W}(\bm{\kappa})=\mathcal{F}_{\mathcal{S}}\{w\}(\bm{\kappa})=\int_{\mathcal{S}}w(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}.$ (74)
@@ IV-B1 Uncoupled CAPA
For uncoupled CAPAs, the optimal beamformer is the matched filtering solution, i.e., $w_{\mathrm{opt}}(\mathbf{s})=\gamma h^{*}(\mathbf{s})$, where $\gamma$ is a scaling factor to satisfy the uncoupled power constraint discussed in Remark 2. Following (III-B2), $\hat{W}(\bm{\kappa})$ for uncoupled CAPAs can be calculated as
$\hat{W}_{\text{uncoupled}}(\bm{\kappa})$ $=\gamma\int_{\mathcal{S}}h^{*}(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s}$
$\hskip-28.45274pt=\gamma\beta_{\mathrm{r}}^{*}L_{x}L_{y}\mathrm{sinc}\left(\frac{\Delta\bar{\kappa}_{x}L_{x}}{2}\right)\mathrm{sinc}\left(\frac{\Delta\bar{\kappa}_{y}L_{y}}{2}\right)$
$\hskip-28.45274pt=\gamma\hat{H}(\bm{\kappa}),$ (75)
where $\Delta\bar{\kappa}_{x}$ and $\Delta\bar{\kappa}_{y}$ are the $x$ and $y$ components of $(\bm{\kappa}-\bm{\kappa}_{\mathrm{r}})$, and $\hat{H}(\bm{\kappa})$ is essentially the windowed 2D Fourier transform of $h^{*}(\mathbf{s})$. Consequently, the uncoupled beampattern is
$B_{\text{uncoupled}}(\bm{\kappa})=\big|\gamma\hat{\beta}(\bm{\kappa})\hat{H}(\bm{\kappa})\big|.$ (76)
@@ IV-B2 Coupled CAPA
Based on Theorem 1, the optimal beamformer for a coupled CAPA is proportional to $v(\mathbf{s})$, i.e., $w_{\mathrm{opt}}(\mathbf{s})=\widetilde{\gamma}v(\mathbf{s})$, where $\widetilde{\gamma}$ is a scaling factor to fulfill the coupled power constraint. The integral equation (28) is a spatial convolution over the finite aperture $\mathcal{S}$. By assuming the aperture is large, the convolution theorem can be applied to approximate this relationship in the wavenumber domain as
$C(\bm{\kappa})\cdot\mathcal{F}_{\mathcal{S}}\{v\}(\bm{\kappa})\approx\hat{H}(\bm{\kappa}),$ (77)
where $C(\bm{\kappa})=Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})$ denotes the Fourier transform of the overall coupling kernel $c(\mathbf{s})$. It is important to note that this approximation is accurate only when the transmit aperture is sufficiently large. For practical finite apertures, this model ignores significant windowing effects. Based on this large-aperture approximation, $\hat{W}(\bm{\kappa})$ for coupled CAPAs is approximated by
$\hat{W}_{\text{coupled}}(\bm{\kappa})=\widetilde{\gamma}\mathcal{F}_{\mathcal{S}}\{v\}(\bm{\kappa})\approx\frac{\widetilde{\gamma}\hat{H}(\bm{\kappa})}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})}.$ (78)
@@
$B_{\text{coupled}}(\bm{\kappa})\approx\left|\frac{\widetilde{\gamma}\hat{\beta}(\bm{\kappa})\hat{H}(\bm{\kappa})}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})}\right|.$ (79)
@@ IV-B3 Impact of Mutual Coupling
$S(\bm{\kappa})=\frac{B_{\text{coupled}}(\bm{\kappa})}{B_{\text{uncoupled}}(\bm{\kappa})}\approx\frac{1}{\varrho\left(Z_{s}+C_{\mathrm{rad}}(\bm{\kappa})\right)},$ (80)
@@ V Extension to Spatially Discrete Arrays
An SPDA model can be obtained from the proposed model of CAPAs. We model the SPDA by partitioning the continuous surface $\mathcal{S}$ into $N$ uniform, non-overlapping antenna elements. Let $\mathcal{S}_{n}\subset\mathcal{S}$ denote the surface of the $n$-th antenna, centered at position $\mathbf{p}_{n}$. We assume that these element surfaces are disjoint, i.e., $\mathcal{S}_{n}\cap\mathcal{S}_{m}=\emptyset$ for $n\neq m$, and each has an identical area $|\mathcal{S}_{n}|=A_{\mathrm{d}}$. Furthermore, we first define a reference element subsurface $\mathcal{S}_{\mathrm{d}}$ centered at the origin as
@@
where $L_{\mathrm{d},x}L_{\mathrm{d},y}=A_{\mathrm{d}}$. The surface of the $n$-th antenna is then a simple translation of this reference, i.e., $\mathcal{S}_{n}=\mathcal{S}_{\mathrm{d}}+\mathbf{p}_{n}$.
@@
where $v_{n}\in\mathbb{C}$ is the discrete beamforming coefficient for the $n$-th antenna, $a_{\mathrm{t}}(\mathbf{s}-\mathbf{p}_{n})$ is the element’s current profile, shifted to its center $\mathbf{p}_{n}$ and $\Pi_{n}(\mathbf{s})$ is a rectangular window function that confines the current to the $n$-th element’s surface, defined as $\Pi_{n}(\mathbf{s})=1$ for $\mathbf{s}\in\mathcal{S}_{n}$ and $\Pi_{n}(\mathbf{s})=0$ otherwise. The CAPA model can be viewed as a continuum idealization of an electrically dense radiating structure. In a practical array, the continuous surface can be realized through finite antenna elements, each occupying a small subsurface and supporting a finite current distribution.
@@ V-A Mutual Coupling
where $\Delta\mathbf{p}_{nm}=\mathbf{p}_{n}-\mathbf{p}_{m}$ denotes the spacing between the $n$-th and the $m$-th antenna elements. This construction makes the connection to practical antenna implementations explicit. Particularly, the discrete coupling matrix is obtained by integrating the continuous kernel over the current profiles of the transmitting elements, which is consistent with the mutual-impedance viewpoint used in conventional array modeling, e.g., [10547020, Eq. (5)]. Moreover, note that the choice of current profile $a_{\mathrm{t}}(\mathbf{s})$ significantly affects the mutual coupling, and it is possible to design $a_{\mathrm{t}}(\mathbf{s})$ carefully to minimize the mutual coupling between different elements.
@@
Based on the analysis in Proposition 1, a half-wavelength spacing cannot null the mutual coupling for polarized antennas, i.e., $c_{\mathrm{rad}}(\Delta\mathbf{p}_{nm})\neq 0$ for $\|\Delta\mathbf{p}_{nm}\|=\lambda/2$. This is in contrast with the case of ideal isotropic antennas that are widely assumed in the literature.
The above SPDA model captures several practical finite-array effects, including finite aperture, finite element area, finite inter-element spacing, and element current profiles. At the same time, it remains a reduced analytical model rather than a geometry-specific full-wave representation of a particular antenna implementation. In particular, effects associated with detailed element shape, feeding structure, and matching networks are not explicitly modeled here. Instead, the model provides a general framework for understanding the fundamental mutual coupling effects in practical arrays and can be used to guide the design of array geometries and element current profiles to mitigate coupling.
@@
Figure 3: Comparison of the true radiation mutual coupling kernel (15) and its approximation in (35) at $7.8$ GHz for different Gauss-Legendre quadrature orders $M$. As $M$ increases, the approximated kernel more accurately reproduces the shape of the true kernel.
@@ V-B Beamforming
where $\bm{h}=[h_{1},\dots,h_{N}]^{H}$ is the channel vector for the SPDA, and $\bm{v}=[v_{1},\dots,v_{N}]^{H}$ is the beamformer vector. Consequently, the beamforming optimization problem for the SPDA can be formulated as
@@ VI Numerical Results
This section presents numerical examples to validate the proposed mutual coupling kernel approximation and the corresponding beamforming designs. Unless specified otherwise, all simulations share the following setup. Specifically, the signal frequency is $f=2.4$ GHz. The transmit surface is a square with dimensions $L_{x}=L_{y}=0.5$ m. The receiver is placed in the far-field at a distance of $R_{0}=50$ m, which satisfies the far-field condition $R_{0}\geq 2D^{2}/\lambda=26$ m, with $D=\sqrt{L_{x}^{2}+L_{y}^{2}}$ denoting the aperture of the transmit surface. The transmit surface is modeled as copper, with conductivity $\sigma_{s}=5.8\times 10^{7}$ S/m and permeability $\mu_{s}=4\pi\times 10^{-7}$ H/m. The Gauss-Legendre order is set to $M=20$. For SPDAs, we assume a simple uniform current profile $a_{\mathrm{t}}(\mathbf{s})=1/\sqrt{A_{\mathrm{d}}}$, an element aperture $0.1\lambda\times 0.1\lambda$ [9906802, wong2014design], and an antenna spacing of $\lambda/2$. Unless otherwise stated, all numerical results correspond to the uni-polarized configuration in Section II-B, where the transmit current is aligned with the $y$-axis and the receiver is polarization matched.
Figure 4: Convergence of the array gain with respect to the Gauss-Legendre quadrature order $M$.
Figure 5: Convergence of the CG method versus the iteration number under different frequencies.
@@
VI-A Convergence of Proposed Methods
We first evaluate the convergence performance of the proposed kernel approximation (KA) and conjugate gradient (CG) methods. First, the accuracy of both methods is highly dependent on the order of the Gauss-Legendre quadrature. In particular, the kernel approximation method exploits the Gauss-Legendre quadrature to approximate the radiation mutual coupling kernel, c.f. (35). Fig. 3 illustrates the true kernel and the approximated kernel under different Gauss-Legendre quadrature orders $M$ at 7.8 GHz. It can be observed that as $M$ increases, the approximated kernel gradually converges to the true kernel, validating the effectiveness of the proposed method. Furthermore, Fig. 5 demonstrates the convergence of the array gain achieved by both methods with respect to $M$. Both methods eventually converge to the same value under different system parameters, which cross-validates the correctness of the results. We note that the CG method converges more smoothly. This may be because KA method exploits the model approximation, i.e., approximating the coupling kernel itself, whereas the CG method applies numerical approximation, i.e., approximating the final numerical implementation. Finally, Fig. 5 shows the convergence of the CG method as iterations proceed. As the frequency increases, the CG method requires more iterations to converge. In contrast, the KA method is non-iterative, thus exhibiting significantly lower computational complexity. Given this advantage, the KA method is used to generate the numerical results in the following sections.
@@ VI-B Array Gain
We now evaluate the array gains under different setups, using both coupled and uncoupled SPDAs as benchmarks. The uncoupled benchmark is obtained by modifying the matrix $\mathbf{\Psi}$ defined in (85), keeping only its diagonal elements and setting all off-diagonal elements to zero.
Fig. 7 depicts the optimal array gain versus the antenna spacing of SPDAs, where smaller spacing implies more antennas are deployed within a fixed aperture. As expected, when mutual coupling is considered, the SPDA array gain gradually converges to the CAPA performance as antenna spacing shrinks. This matches the physical expectation that CAPA is the ultimate limit of the SPDA. Nevertheless, the uncoupled model fails to capture the correct behavior. It underestimates performance at large spacing and significantly overestimates it at small spacing. Specifically, the array gain for the uncoupled SPDAs tends to increase unboundedly as spacing reduces, which is a violation of physics. Fig. 7 further shows the array gain versus the aperture size $|\mathcal{S}|$. For coupled CAPAs, the array gain grows linearly as the aperture size increases. However, the SPDA array gain exhibits a stepwise increase, since a larger aperture does not always accommodate more antenna elements.
@@ VII Conclusions
This paper has primarily focused on uni-polarized and single-user CAPA systems. Although the underlying mutual coupling model is developed in a general tri-polarized form, the explicit beamforming design and numerical evaluation have been specialized to a fixed uni-polarized setting for analytical tractability. Several promising future research directions can be identified as follows:
@@
  • Full-Wave and Experimental Validation: While this work develops a physics-based analytical model and validates its internal predictions numerically, an important next step is to benchmark the model against geometry-specific full-wave simulations and experimental measurements for finite practical arrays, including implementation-level effects such as feeding and matching constraints.